6. (Sprcadshect Problem) You are given that the random variable X is distributed as a Gcometric distribution with paramcters B-3. Calculate: a. Pr[X5]
Answer #1

For Geometric distribution, we have PDF as

1 3 3

a)

P(X-5) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4) + P(X = 5) 1 21 41 8 16 1 321 3 3 39 3 27 381 3 243 3 1 243+2(81) 4(27) +8(9) 16(3) + 32 243 0.9122

b)

8 116 1 32 1 27 3 81 3 243 3 1 (8(9) +16(3) 32 0 3801 243 0.2085